OpenStudy (anonymous):

Extra Credit Problems? Will fan and medal! 1: Factor completely 5x(x + 3) + 6(x + 3). A. (x + 3)(30x) B. (x + 3)(5x + 6) C. (x + 3)(11x) D. prime 2: Factor completely 4ab − 7ax + 8b − 14x. A. (4b − 7x)(a + 2) B. (4b + 7x)(a + 2) C. (4b − 7x)(a − 2) D.prime 3: Factor completely 9xy + 36x − 10y − 40. A. prime B. (y − 4)(9x − 10) C. (y + 4)(9x + 10) D. (y + 4)(9x − 10)

OpenStudy (campbell_st):

I won't answer them, that's your task... I won't get the extra credit... you will but I can explain how it works... in the 1st question what is the obvious common factor...?

OpenStudy (owlcoffee):

On the first one, Try doing the common factor, with that, I mean the factor that is repeated on both terms. As an example: \[a^2(l+5+b^2)-b^2(l+5+b^2)-(l+5+b^2)\] If I multiply and divide by (l+5+b^2) I get this: \[\frac{ (l+5+b^2)(a^2(l+5+b^2)-b^2(l+5+b^2)-(l+5+b^2)) }{ (l+5+b^2) }\] And doing some very simple algebra: \[(l+5+b^2)(\frac{ a^2(l+5+b^2) }{ (l+5+b^2) }-\frac{ b^2(l+5+b^2) }{ (l+5+b^2) }+\frac{ (l+5+b^2) }{ (l+5+b^2) })\] And operating the denominators with the numerators: \[(l+5+b^2)(a^2-b^2+1)\] That would be The "long way" to apply the "common factor" axiom.